# Rethinking External Slow-Thinking: From Snowball Errors to Probability of Correct Reasoning

### v2

@@ 3 From Snowball Errors to Probability
Figure 2: The esimated MI and reward of the responses generated by different LLMs on GSM8k. The x-axis denotes the average length of the responses. The left y-axis represents the estimated MI between each response and the golden answer, whereas the right y-axis shows the corresponding reward. $\bullet$” indicates the estimated MI for individual questions, while “ ” and “ ” respectively depict the fitted curves of the estimated MI and reward. “ ” depicts the linear baseline.
@@ 3.1 Probability of Reasoning Errors
The proof mainly relies on Fano’s inequality (Fano, 2008) and is provided in Section A.2. Theorem 3.3 sets a lower bound on the probability of reasoning errors at step $l$. Since $H_{<l}(\bm{t}|\bm{r})$ represents the cumulative information loss up to step $l$, it is assumed to increase at least linearly with $l$. Furthermore, when the snowball effect occurs, $H_{<l}(\bm{t}|\bm{r})$ may grow faster than linearly. As a result, the lower bound on the probability of reasoning errors rises with the reasoning path length $l$, indicating that the chance of errors increases as snowball errors accumulate.
@@ 3.2 Empirical Verification
To empirically verify the existence of snowball errors in LLM reasoning, we conducted experiments using the GSM8k dataset (Cobbe et al., 2021). The evaluation was performed on three state-of-the-art reasoning LLMs: Llama-3.1-8B-Instruct (2024), Qwen2.5-7B-Instruct (2024a), and Skywork-o1-Open-Llama-3.1-8B (2024). We estimate the mutual information $I(\bm{t};\bm{r})$ per token and the reward of the responses to illustrate their relationship with the reasoning path length $L$. Detailed experimental settings are provided in Section B.1.
As illustrated in Figure 2, the estimated mutual information decreases in a nearly exponential and much faster rate than linear decay as $L$ increases. Since our verification computes the average mutual information per token, the actual rate of decline in mutual information for tokens appearing later in the sequence is likely to be more pronounced. Furthermore, the reward scores of the responses correspond to the mutual information, and also diminish with increasing response length. These findings confirm the existence of snowball errors by verifying that the information loss can accumulate much faster than linear in the reasoning processes of LLMs, and illustrate their relevance to the quality of responses, aligning with our theoretical analysis.
@@ 4.2 Probability of Correct Reasoning
Furthermore, we can derive the probability of generating a $\tau$-correct response $\mathcal{R}$ as follows:
@@ 4.3 External Slow-Thinking Mechanisms
The proof is provided in Section A.4. Lemma 4.5 and Theorem 4.6 show that the probability of generating a correct response in width-expanding methods depends on three key factors: the number of generations $k$ per layer, the number of candidates $b$ selected per layer, and the correctness threshold $\tau$. Specifically, components related to $\epsilon_{b}$ represent $\operatorname{Pr}\left(\tau_{\text{select}}\right)$, which are influenced by the reliability of the value function used in selection. Meanwhile, $\operatorname{Pr}\left(\tau_{\text{generate}}\right)$ is mainly determined by the expansion width $k$ and reasoning path length $L$. Compared with Lemma 4.4, Theorem 4.6 further illustrates the mechanism of external slow-thinking methods. By scaling $k$, these methods can improve the probability of correct reasoning at the cost of additional reasoning steps. However, the effectiveness heavily depends on the reliability of the value function, reflected in $\epsilon_{b}^{L}$. For slow-thinking methods to be effective, the value function must satisfy $\epsilon_{b}>\frac{1}{k}$ to ensure an improvement in the probability upper bounds.
Based on the above analysis, we can infer the underlying mechanisms of external slow-thinking methods. By expanding the reasoning space, these methods effectively increase the probability of generating a correct response, thereby mitigating the impact of snowball errors. However, selecting the most promising reasoning path poses a significant challenge, as the effectiveness of this selection heavily depends on the reliability of the employed value function, which can substantially influence the overall performance of the method. We also provide analyses of the mechanism of external slow-thinking methods from other perspective in Appendix C.
@@ 5.3 Empirical Evaluation
We empirically compare the accuracy of BoN and MCTS on two reasoning tasks: GSM8k (Cobbe et al., 2021) and PrOntoQA (Saparov & He, 2022). Following prior work (Feng et al., 2023), we use the recommended settings to optimize MCTS and calculate the corresponding $N$ for BoN to align its reasoning cost with MCTS. Due to differences in how reasoning paths are generated, exact alignment is impractical. To address this, we define $N$ as reasonable if it lies between $\widetilde{N}_{\text{res}}$ (aligned reasoning steps) and $\widetilde{N}_{\text{call}}$ (aligned LLM calls). We evaluate BoN with three selection strategies: Self-Consistency, ORM Vote, and ORM Max. Details of the experimental setup are in Section B.2, and results are presented in Figure 3.
@@ 6 Related Work
Reasoning with LLMs. LLMs have made substantial progress in understanding and generation, particularly for complex reasoning tasks. Foundational work (Brown et al., 2020) established LLMs as few-shot learners, while Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduced multi-step reasoning by explicitly generating intermediate steps. Self-Consistency (Wang et al., 2022) further improved robustness by aggregating multiple reasoning paths. Information-theoretic approaches, such as (Ton et al., 2024), have also provided theoretical insights into quantifying and mitigating reasoning errors, showcasing the interplay between empirical advancements and theoretical frameworks.
External Slow-Thinking. Test-time scaling has proven effective in enhancing LLM reasoning (Snell et al., 2024). External slow-thinking methods, such as generating additional tokens (Wei et al., 2022) and incorporating tree search algorithms (Yao et al., 2024), expand the reasoning space without retraining. Techniques like beam search (Kang et al., 2024) and Monte Carlo Tree Search (MCTS) (Zhang et al., 2024; Feng et al., 2023) have further advanced reasoning capabilities. However, these methods require significantly more computational resources compared to simpler strategies like Best-of-N and often yield limited practical improvement. The underlying mechanisms of their effectiveness remain an area of ongoing exploration.
@@ Appendix A Proofs
Appendix B Experiment Settings
B.1 Settings for Verifying Snowball Errors
@@
B.2 Settings for Comparison between BoN and MCTS
@@
In our experiments, we compared the performance of BoN and MCTS in the context of the Snowball task. We first determine the baseline MCTS setting according to the recommendation in previous work (Feng et al., 2023). Specifically, in GSM8k, we set the tree max width with $6$ Tree Max depth with $8$. While in PrOntoQA, we set the tree max width with $6$ and Tree Max depth with $15$.
@@
GSM8k PrOntoQA
avg. $b$ 4.26 1.67
avg. $p$ 4.54 9.45
avg. $L$ 3.11 4.00
$\widetilde{N}_{\text{call}}$ 19.40 15.77
$\widetilde{N}_{\text{res}}$ 6.23 3.94
Table 2: Settings for BoN and MCTS in GSM8k and PrOntoQA.
Appendix C More External Slow-Thinking Mechanisms Analysis
@@
With theorem C.4, we can derive the best setting of rollout step $\gamma$ in Lookahead Search.
Corollary C.5.

### v3

@@ 3 From Snowball Errors to Probability
Figure 2: The estimated MI and reward of the responses generated by different LLMs on GSM8k. The x-axis denotes the average length of the responses. The left y-axis represents the estimated MI between each response and the golden answer, whereas the right y-axis shows the corresponding reward. $\bullet$” indicates the estimated MI for individual questions, while “ ” and “ ” respectively depict the fitted curves of the estimated MI and reward. “ ” depicts the linear baseline.
@@ 3.1 Probability of Reasoning Errors
The proof mainly relies on Fano’s inequality (Fano, 2008) and is provided in Section A.2. Theorem 3.3 sets a lower bound on the probability of reasoning errors at every intermediate step $l$. Since $H_{<l}(\bm{t}|\bm{r})$ represents the cumulative information loss up to step $l$, it is assumed to increase at least linearly with $l$. Furthermore, when the snowball effect occurs, $H_{<l}(\bm{t}|\bm{r})$ may grow faster than linearly. As a result, the lower bound on the probability of reasoning errors rises with the reasoning path length $l$, indicating that the chance of errors increases as snowball errors accumulate.
@@ 3.2 Empirical Verification
To empirically verify the existence of snowball errors in LLM reasoning, we conducted experiments using the GSM8k dataset (Cobbe et al., 2021). The evaluation was performed on three state-of-the-art reasoning LLMs: Llama-3.1-8B-Instruct (2024), Qwen2.5-7B-Instruct (2024a), and Skywork-o1-Open-Llama-3.1-8B (2024). We estimate the mutual information $I(\bm{t};\bm{r})$ per token and the reward of the responses to illustrate their relationship with the reasoning path length $L$. Detailed experimental settings are provided in Section C.1.
As illustrated in Figure 2, the estimated mutual information decreases in a nearly exponential and much faster rate than linear decay as $L$ increases. Since our verification computes the average mutual information per token, the actual rate of decline in mutual information for tokens appearing later in the sequence is likely to be more pronounced. Furthermore, the reward scores of the responses correspond to the mutual information, and also diminish with increasing response length. These findings confirm the existence of snowball errors by verifying that the information loss can accumulate much faster than linearly in the reasoning processes of LLMs, and illustrate their relevance to the quality of responses, aligning with our theoretical analysis. For a more robust verification, we also include results on larger LLMs for a more robust verification in Section D.1 and an additional analysis by different difficulty levels in Section D.2.
@@ 4.2 Probability of Correct Reasoning
This proposition mainly aims to obtain more intuitive conclusions in subsequent analysis and is based on the intuition of LLM’s multi-step reasoning errors. In particular, we point out in Appendix B that under a more relaxed and generalized proposition, our subsequent theoretical results can still be guaranteed.
With Proposition 3.1, we can derive the probability of generating a $\tau$-correct response $\mathcal{R}$ as follows:
@@ 4.3 External Slow-Thinking Mechanisms
The proof is provided in Section A.4, and we also provide an empirical analysis about the influence of $k$ in Section D.3. Lemma 4.5 and Theorem 4.6 show that the probability of generating a correct response in width-expanding methods depends on three key factors: the number of generations $k$ per layer, the number of candidates $b$ selected per layer, and the correctness threshold $\tau$. Specifically, components related to $\epsilon_{b}$ represent $\operatorname{Pr}\left(\tau_{\text{select}}\right)$, which are influenced by the reliability of the value function used in selection. Meanwhile, $\operatorname{Pr}\left(\tau_{\text{generate}}\right)$ is mainly determined by the expansion width $k$ and reasoning path length $L$. Compared with Lemma 4.4, Theorem 4.6 further illustrates the mechanism of external slow-thinking methods. By scaling $k$, these methods can improve the probability of correct reasoning at the cost of additional reasoning steps. However, the effectiveness heavily depends on the reliability of the value function, reflected in $\epsilon_{b}^{L}$. For slow-thinking methods to be effective, the value function must satisfy $\epsilon_{b}>\frac{1}{k}$ to ensure an improvement in the probability upper bounds.
Based on the above analysis, we can infer the underlying mechanisms of external slow-thinking methods. By expanding the reasoning space, these methods effectively increase the probability of generating a correct response, thereby mitigating the impact of snowball errors. However, selecting the most promising reasoning path poses a significant challenge, as the effectiveness of this selection heavily depends on the reliability of the employed value function, which can substantially influence the overall performance of the method. We also provide analyses of the mechanism of external slow-thinking methods from other perspective in Appendix E.
@@ 5.3 Empirical Evaluation
We empirically compare the accuracy of BoN and MCTS on two reasoning tasks: GSM8k (Cobbe et al., 2021) and PrOntoQA (Saparov & He, 2022). Following prior work (Feng et al., 2023), we use the recommended settings to optimize MCTS and calculate the corresponding $N$ for BoN to align its reasoning cost with MCTS. Due to differences in how reasoning paths are generated, exact alignment is impractical. To address this, we define $N$ as reasonable if it lies between $\widetilde{N}_{\text{res}}$ (aligned reasoning steps) and $\widetilde{N}_{\text{call}}$ (aligned LLM calls). We evaluate BoN with three selection strategies: Self-Consistency, ORM Vote, and ORM Max. Details of the experimental setup are in Section C.2, and results are presented in Figure 3. Furthermore, we also include additional experimental verifications on a planning task Game24 (Yao et al., 2024) in Section D.4.
@@ 6 Related Work
Reasoning with LLMs. LLMs have made substantial progress in understanding and generation, particularly for complex reasoning tasks. Foundational work (Brown et al., 2020) established LLMs as few-shot learners, while Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduced multi-step reasoning by explicitly generating intermediate steps. Self-Consistency (Wang et al., 2022) further improved robustness by aggregating multiple reasoning paths. Information-theoretic approaches  (Ton et al., 2024) have also provided theoretical insights into quantifying and mitigating reasoning errors, showcasing the interplay between empirical advancements and theoretical frameworks.
External Slow-Thinking. Test-time scaling has proven effective in enhancing LLM reasoning (Snell et al., 2024), this timely inference scaling law has been empirically analyzed and verified (Wu et al., 2024). External slow-thinking methods, such as generating additional tokens (Wei et al., 2022) and incorporating tree search algorithms (Yao et al., 2024), expand the reasoning space without retraining. Techniques like beam search (Kang et al., 2024) and Monte Carlo Tree Search (MCTS) (Zhang et al., 2024; Feng et al., 2023) have further advanced reasoning capabilities. However, these methods require significantly more computational resources compared to simpler strategies like Best-of-N and often yield limited practical improvement. The underlying mechanisms of their effectiveness remain an area of ongoing exploration.
@@ 7 Conclusion
Acknowledgments
This research was supported by National Natural Science Foundation of China (No.62476277), National Key Research and Development Program of China (No. 2024YFE0203200), CCF-ALIMAMA TECH Kangaroo Fund (No.CCF-ALIMAMA OF 2024008), and Huawei-Renmin University joint program on Information Retrieval. We also acknowledge the support provided by the fund for building worldclass universities (disciplines) of Renmin University of China and by the funds from Beijing Key Laboratory of Big Data Management and Analysis Methods, Gaoling School of Artificial Intelligence, Renmin University of China, from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “DoubleFirst Class” Initiative, Renmin University of China, from Public Policy and Decision-making Research Lab of Renmin University of China, and from Public Computing Cloud, Renmin University of China.
Impact Statement
This paper presents work whose goal is to advance the field of Large Language Models. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
Appendix B Relaxed Theoretical Results
Proposition 4.3 was designed to facilitate subsequent analyses while maintaining readability. The negative exponential form provides a simple yet effective characterization of accuracy decay. Below, we would like to prove that our main results remain valid under a weaker assumption when $\operatorname{Pr}\left[|\phi(r_{l})-\phi(r_{l}^{*})|\leq\tau\right]$ decreases monotonically with $l$.
Proposition B.1.
(Relaxed probability of $\tau$-correct step for Proposition 4.3.) Instead of requiring $\operatorname{Pr}\left[|\phi(r_{l})-\phi(r_{l}^{*})|\leq\tau\right]=\operatorname{min}\left(\lambda_{\tau}e^{-l},1\right)$, we now assume only that the left-hand side decreases monotonically with $l$ and converges to 00, i.e.,
$\operatorname{Pr}\left[|\phi(r_{l})-\phi(r_{l}^{*})|\leq\tau\right]=\operatorname{min}\left(\xi(l,\tau),1\right),$
where $\xi(l,\tau)\geq 0$ decreases monotonically with $l$ and converges to 00.
Under this weaker condition, by noting that $\xi^{L}(l,\tau):=\prod_{l=1}^{L}\xi(l,\tau)$, we derive revised relaxed bounds for subsequent results in the main text presented below:
Lemma B.2.
(Relaxed probability of $\tau$-correct reasoning for Lemma 4.4.) After relaxation, the probability of generating a $\tau$-correct response $\mathcal{R}$ is:
$\leq\xi^{L}(l,\tau).$
Lemma B.3.
(Relaxed probability of $\tau$-correct reasoning in width-expanding methods for Lemma 4.5.) After relaxation, for a beam-search-like strategy which samples $k$ steps and keeps $b$ steps as candidates at each expansion, the probability of obtaining a $\tau$-correct response is upper bounded by:
$\operatorname{Pr}\left[\psi(\mathcal{R})\leq\tau\right]\leq\prod_{l=1}^{L}\epsilon_{b}\left[1-\left(1-\xi(l,\tau)\right)^{k}\right].$
Theorem B.4.
(Relaxed upper bound of the probability of $\tau$-correct reasoning in width-expanding methods for Theorem 4.6.) After relaxation, the probability of obtaining a $\tau$-correct response is upper bounded by:
$\operatorname{Pr}\left[\psi(\mathcal{R})\leq\tau\right]\leq\epsilon_{b}^{L}k^{L}\xi^{L}(l,\tau).$
Lemma B.5.
(Relaxed upper bound of the probability of $\tau$-correct reasoning with BoN for Lemma 5.1.) After relaxation, the probability of obtaining a $\tau$-correct response in BoN satisfies:
$\operatorname{Pr}\left[\psi(\mathcal{R})\leq\tau\right]\leq\epsilon_{N}N^{L}\xi^{L}(l,\tau).$
Lemma B.6.
(Relaxed upper bound of the probability of $\tau$-correct reasoning with MCTS in best case for Lemma 5.2.) After relaxation, the probability of obtaining a $\tau$-correct response in MCTS in the best case satisfies:
$\operatorname{Pr}\left[\psi(\mathcal{R})\leq\tau\right]\leq\epsilon_{b}^{L}b^{L}\xi^{L}(l,\tau).$
Lemma B.7.
(Relaxed upper bound of the probability of $\tau$-correct reasoning with MCTS in worst case for Lemma 5.3.) After relaxation, the probability of generating a $\tau$-correct response in MCTS in the worst case satisfies:
$\operatorname{Pr}\left[\psi(\mathcal{R})\leq\tau\right]\leq b^{\frac{L(L+1)}{2}}\xi^{L}(l,\tau)\prod_{l=1}^{L}\epsilon_{b^{l}}.$
These modifications of relaxed results preserve the validity of Corollary 5.4, Corollary 5.5, and Table 1, confirming that our theoretical results are robust even without the original exponential form in Proposition 3.1.
Appendix C Experiment Settings
C.1 Settings for Verifying Snowball Errors
@@
C.2 Settings for Comparison between BoN and MCTS
@@
In our experiments, we compared the performance of BoN and MCTS in the context of the Snowball task. We first determine the baseline MCTS setting according to the recommendation in previous work (Feng et al., 2023). Specifically, in GSM8k, we set the tree max width to $6$ Tree Max depth to $8$. In PrOntoQA, we set the tree max width to $6$ and Tree Max depth to $15$. And in Game24, we set the tree max width to $20$ and Tree Max depth to $4$.
@@
GSM8k PrOntoQA Game24
avg. $b$ 4.26 1.67 4.56
avg. $p$ 4.54 9.45 3.99
avg. $L$ 3.11 4.00 3.00
$\widetilde{N}_{\text{call}}$ 19.40 15.77 18.24
$\widetilde{N}_{\text{res}}$ 6.23 3.94 6.08
Table 2: Settings for BoN and MCTS in GSM8k, PrOntoQA and Game24.
Appendix D Additional Empirical Verification
D.1 Snowball Error Verifications on Larger LLMs
For a more robust verification of our theoretical findings, we have conducted extensive analyses on larger language models (Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct) in addition to the 7B/8B models presented in Figure 2. We maintained identical experimental settings and workflow as described in the original experiments to ensure methodological consistency, and we present the additional results in Figure 4.
Our key findings from these additional experiments demonstrate that the MI decay pattern remains consistent across larger model sizes, exhibiting similar behavior to smaller models. Also, response quality continues to show a negative correlation with output length, as observed in our original experiments.
Figure 4: Additional snowball error verifications. We have conducted extra experiments on larger LLMs, including Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct, with the same setting as results in Figure 2 of our main text.
D.2 Snowball Error Analysis by Difficulty Levels
We have conducted verifications for the snowball errors at different difficulty levels. The experiments are performed on MATH-500 (Lightman et al., 2023), where the questions are divided into 5 different difficulty levels (level 1 for the simplest, level 5 for the most difficult). The results are presented in Figure 5. In (a), we present the relationship between the estimated MI and average response length at different levels. For all 5 levels, we observe similar MI decay phenomena like Figure 2 in our main text. In (b), we present the relationship between the average accuracy and the length of responses at different levels. For all 5 levels, we observed that accuracy generally decreases with response length. Furthermore, the accuracy of the simplest level 1 decreases significantly when the length increases, demonstrating the harm of overthinking. These results demonstrate that the MI decay phenomenon is highly relevant to the response length even when the influence of question difficulty is similar.
Figure 5: Snowball error analysis by different difficulty levels. We have conducted extra experiments to analyze the mutual information (MI) and accuracy at different difficulty levels. The questions at the same level share similar difficulty. (a) The MI analysis by difficulty levels. (b) The accuracy analysis by different difficulty levels.
D.3 Influence of $k$
For better understanding Theorem 4.6, we have conducted an analysis about the influence of $k$ in the theoretical findings, which is presented in Figure 6. The experiments are performed on GSM8k and PrOntoQA.
The results illustrate that: (1) the reasoning correctness increases when the reasoning cost $k$ increases, which aligns with the theoretical findings in Theorem 4.6; (2) for simpler PrOntoQA task where the value function is more reliable, the accuracy increases faster than GSM8k when $k$ increases.
Figure 6: Influence of $k$. We have conducted extra experiments to further verify Theorem 4.6 by examining the influence of $k$. The experiments are performed on GSM8k and PrOntoQA, and we vary the value of the max-action of an MCTS while keeping other hyperparameters optimal, according to previous studies.
D.4 Additional Verification for BoN v.s. MCTS
We conducted additional verification experiments using the Game24 benchmark (Yao et al., 2024), which requires solving arithmetic puzzles to obtain the number 24. The setting is similar with Figure 3, and we present the results in Table 3.
This result illustrates a similar conclusion to Fig.3, when the total reasoning cost is comparable, BoN can achieve a comparable and even better performance than MCTS.
$N$ 2 4 6 8 10 12 14 16 18 20
ORM-Vote 17.96% 29.01% 38.67% 46.13% 50.83% 53.87% 59.94% 63.81% 66.85% 67.68%
ORM-Max 17.96% 29.01% 38.67% 46.13% 50.83% 53.87% 59.94% 63.81% 67.13% 68.23%
Table 3: Additional verification for BoN v.s. MCTS in Game24. The baseline MCTS obtain an accuracy of 64.80% where $\widetilde{N}_{\text{res}}=6.08$ and $\widetilde{N}_{\text{call}}=18.24$.
Appendix E More External Slow-Thinking Mechanisms Analysis
@@
With theorem E.4, we can derive the best setting of rollout step $\gamma$ in Lookahead Search.
Corollary E.5.